# Identifying Trustworthiness Challenges in Deep Learning Models for Continental-Scale Water Quality Prediction

### v1

@@ Acknowledgements
| USGS code | Description | Abbreviation | Unit | # Observations per basin |
| 00010 | Water temperature | Temp | °C | 330.5 |
| 00095 | Specific conductance | Cond | uS/cm at 25°C | 285.6 |
| 00300 | Oxygen | DO | mg/L | 197.8 |
| 00040 | pH | pH | - | 224.9 |
| 00405 | Carbon dioxide | mg/L | 129.2 | |
| 00600 | Total nitrogen | TN | mg/L | 193.3 |
| 00605 | Organic nitrogen | OrgN | mg/L | 171.7 |
| 00618 | Nitrate | mg/L as N | 138.3 | |
| 00660 | Orthophosphate | mg/L as | 204.9 | |
| 00665 | Total phosphorus | TP | mg/L as P | 266.9 |
| 00681 | Organic carbon | NPOC | mg/L | 60.3 |
| 00915 | Calcium | mg/L | 131.7 | |
| 00925 | Magnesium | mg/L | 131.8 | |
| 00930 | Sodium | mg/L | 117.3 | |
| 00935 | Potassium | mg/L | 114.8 | |
| 00940 | Chloride | mg/L | 184.1 | |
| 00945 | Sulfate | mg/L | 154.3 | |
| 00955 | Silica | mg/L | 116.1 | |
| 71846 | Ammonia and ammonium | mg/L as | 184.1 | |
| 80154 | Suspended sediment concentration | TSS | mg/L | 305.4 |
| Group | Name | Type | Description | Unit |
| Runoff | runoff | time-varying | Area normalized streamflow from USGS | m/y |
| Meteorological forcings | pr | time-varying | Daily total precipitation | mm/day |
| sph | time-varying | Specific humidity | ||
| srad | time-varying | Surface downwelling solar radiation | ||
| tmmn | time-varying | Daily minimum 2-meter air temperature | F | |
| tmmx | time-varying | Daily maximum 2-meter air temperature | F | |
| pet | time-varying | Reference grass evapotranspiration | mm/day | |
| etr | time-varying | Reference alfalfa evapotranspiration | mm/day | |
| Rainfall chemistry | pH | time-varying | Logarithm of the H ion activity | unitless |
| Cond | time-varying | Electrical conductivity of water | ||
| time-varying | Ca ion concentration | mg/L | ||
| time-varying | Mg ion concentration | mg/L | ||
| time-varying | K ion concentration | mg/L | ||
| time-varying | Na ion concentration | mg/L | ||
| time-varying | concentration | mg/L | ||
| time-varying | concentration | mg/L | ||
| time-varying | Cl ion concentration | mg/L | ||
| time-varying | concentration | mg/L | ||
| distNTN | time-varying | The distance to the nearest NTN sampling site | km | |
| Vegetation indices | LAI | time-varying | Leaf area index of vegetation | |
| FAPAR | time-varying | Fraction of absorbed photosynthetically active radiation | unitless | |
| NPP | time-varying | Net primary production | ||
| Time variables | datenum | time-varying | The number of days relative to January 1, 2000 | unitless |
| sinT | time-varying | Sine of datenum | unitless | |
| cosT | time-varying | Cosine of datenum | unitless | |
| Basic characteristics | HYDRO_DISTURB_INDX | static | Hydrologic “disturbance index” score, based on 7 variables: 1) MAJ_DDENS_2009, 2) WATER_WITHDR, 3) change in dam storage 1950-2009, 4) CANALS_PCT, 5) RAW_DIS_NEAREST_MAJ_NPDES, 6) ROADS_KM_SQ_KM, and 7) FRAGUN_BASIN | unitless |
| BAS_COMPACTNESS | static | Watershed compactness ratio, = area/ * 100; higher number = more compact shape | unitless | |
| DRAIN_SQKM | static | Watershed drainage area, sq km, as delineated in our basin boundary | ||
| Geology | GEOL_REEDBUSH_DOM | static | Dominant (highest percent of area) geology, derived from a simplified version of Reed & Bush (2001) - Generalized Geologic Map of the Conterminous United States | unitless |
| GEOL_REEDBUSH_DOM_PCT | static | Percentage of the watershed covered by the dominant geology type | percentage | |
| Hydrologic characteristics | STREAMS_K_S_KM | static | Stream density, km of streams per watershed sq km, from NHD 100k streams | |
| STRAHLER_MAX | static | Maximum Strahler stream order in the watershed, from NHDPlus | unitless | |
| MAINSTEM_SINUOUSITY | static | Sinuosity of mainstem stream line, from our delineation of mainstem stream lines. Defined as curvilinear length of the mainstem stream line dividedby the straight-line distance between the end points of the line. | unitless | |
| BFI_AVE | static | Base Flow Index (BFI). The BFI is a ratio of base flow to total streamflow, expressed as a percentage and ranging from 0 to 100. Base flow is the sustained, slowly varying component of streamflow, usually attributed to ground-water discharge to a stream. | percentage | |
| CONTACT | static | Subsurface flow contact time index | days | |
| PCT_1ST_ORDER | static | Percent of stream lengths in the watershed which are first-order streams (Strahler order); from NHDPlus & percentage | percentage | |
| PCT_2ND_ORDER | static | Percent of stream lengths in the watershed which are second-order streams (Strahler order); from NHDPlus & percentage | percentage | |
| PCT_3RD_ORDER | static | Percent of stream lengths in the watershed which are third-order streams (Strahler order); from NHDPlus & percentage | percentage | |
| PCT_4TH_ORDER | static | Percent of stream lengths in the watershed which are fourth-order streams (Strahler order); from NHDPlus & percentage | percentage | |
| PCT_5TH_ORDER | static | Percent of stream lengths in the watershed which are fifth-order streams (Strahler order); from NHDPlus & percentage | percentage | |
| PCT_6TH_ORDER_OR_MORE | static | Percent of stream lengths in the watershed which are sixth or greater-order streams (Strahler order); from NHDPlus & percentage | percentage |
| Group | Name | Type | Description | Unit |
| Historical and current dams information | DDENS_2009 | static | Dam density; number per 100 km sq | number of dams/100 |
| STOR_NOR_2009 | static | Dam storage in watershed (“NORMAL_STORAGE”); megaliters total storage per sq km (1 megalitres = 1,000,000 liters = 1,000 cubic meters) | megaliters/ | |
| NPDES | NPDES_MAJ_DENS | static | Density of NPDES (National Pollutant Discharge Elimination System) “major” point locations in the watershed; number per 100 km sq. Major locations are defined by an EPA-assigned major flag. From the download of NPDES national database summer 2006. | number of sites/100 |
| Percentages of land cover 2006 in the watershed and lanscape | DEVNLCD06 | static | Watershed percent “developed” (urban), 2006 era (2001 for AK-HI-PR). Sum of classes 21, 22, 23, and 24. | percentage |
| FORESTNLCD06 | static | Watershed percent “forest”, 2006 era (2001 for AK-HI-PR). Sum of classes 41, 42, and 43. | percentage | |
| PLANTNLCD06 | static | Watershed percent “planted/cultivated” (agriculture), 2006 era (2001 for AK-HI-PR). Sum of classes 81 and 82. | percentage | |
| WATERNLCD06 | static | Watershed percent Open Water (class 11) | percentage | |
| WOODYWETNLCD06 | static | Watershed percent Woody Wetlands (class 90) | percentage | |
| EMERGWETNLCD06 | static | Watershed percent Emergent Herbaceous Wetlands (class 95) | percentage | |
| Nitrogen and phosphorus application rate in the watershed | NITR_APP_KG_SQKM | static | Estimate of nitrogen from fertilizer and manure, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed. | kg/ |
| PHOS_APP_KG_SQKM | static | Estimate of nitrogen from fertilizer and manure, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed. | kg/ | |
| Pesticide | PESTAPP_KG_SQKM | static | Estimate of agricultural pesticide application (219 types), kg/sq km, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed | kg/ |
| Regions | ECO2_BAS_DOM | static | Dominant (highest % of the area) Level II ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. | unitless |
| ECO3_BAS_DOM | static | Dominant (highest % of the area) Level III ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. | Level III ecoregion (1-84) | |
| NUTR_BAS_DOM | static | Dominant (highest % of the area) nutrient ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. | Nutrient ecoregion (1-14) | |
| HLR_BAS_DOM_100M | static | Dominant (highest % of the area) Hydrologic Landscape Region within the watershed. See X_Region_Names sheet for crosswalk to name. | HLR region (1-20) | |
| PNV_BAS_DOM | static | Dominant (highest % of the area) Potential Natural Vegetation (PNV) within the watershed. See X_Region_Names sheet for crosswalk to name. | PNV type (1-63) | |
| Soil | AWCAVE | static | Average value for the range of available water capacity for the soil layer or horizon (inches of water per inch of soil depth) | unitless |
| PERMAVE | static | Average permeability (inches/hour) | inches/hour | |
| BDAVE | static | Average value of bulk density (grams per cubic centimeter) | grams per cubic centimeter | |
| OMAVE | static | Average value of organic matter content (percent by weight) | percentage | |
| WTDEPAVE | static | Average value of depth to seasonally high water table (feet) | feet | |
| ROCKDEPAVE | static | Average value of total soil thickness examined (inches) | inches | |
| CLAYAVE | static | Average value of clay content (percentage) | percentage | |
| SILTAVE | static | Average value of silt content (percentage) | percentage | |
| KFACT_UP | static | Average K-factor value for the uppermost soil horizon in each soil component. K-factor is an erodibility factor which quantifies the susceptibility of soil particles to detachment and movement by water. The K-factor is used in the Universal Soil Loss Equation (USLE) to estimate soil loss by water. Higher values of the K-factor indicate greater potential for erosion | unitless | |
| RFACT | static | Rainfall and Runoff factor (“R factor” of Universal Soil Loss Equation); average annual value for the period 1971-2000. | 100s ft-tonf in/h/ac/yr | |
| Topographic characteristics | ELEV_MEAN_M_BASIN | static | Mean watershed elevation (meters) from 100m National Elevation Dataset | m |
| SLOPE_PCT | static | Mean watershed slope, percent. Derived from 100m resolution National Elevation Dataset, so slope values may differ from those calculated from data of other resolutions. | percentage | |
| ASPECT_DEGREES | static | Mean watershed aspect, degrees (degrees of the compass, 0-360). Derived from 100m resolution National Elevation Data. 0 and 360 point to north, because of the national Albers projection actual aspect may vary. | degrees (0-360) | |
| Latitude and Longitude | LAT_GAGE | static | Latitude at gage, decimal degrees | decimal degrees, datum NAD83 |
| LNG_GAGE | static | Longitude at gage, decimal degrees | decimal degrees, datum NAD83 | |
| Snow | SNOW_PCT_PRECIP | static | Snow percent of total precipitation estimate, mean for period 1901-2000. From McCabe and Wolock (submitted, 2008), 1km grid. | percentage |

### v2

@@ Acknowledgements
USGS code Description Abbreviation Unit # Observations per basin 00010 Water temperature Temp °C 330.5 00095 Specific conductance Cond uS/cm at 25°C 285.6 00300 Oxygen DO mg/L 197.8 00400 pH pH - 224.9 00405 Carbon dioxide mg/L 129.2 00600 Total nitrogen TN mg/L 193.3 00605 Organic nitrogen OrgN mg/L 171.7 00618 Nitrate mg/L as N 138.3 00660 Orthophosphate mg/L as 204.9 00665 Total phosphorus TP mg/L as P 266.9 00681 Organic carbon NPOC mg/L 60.3 00915 Calcium mg/L 131.7 00925 Magnesium mg/L 131.8 00930 Sodium mg/L 117.3 00935 Potassium mg/L 114.8 00940 Chloride mg/L 184.1 00945 Sulfate mg/L 154.3 00955 Silica mg/L 116.1 71846 Ammonia and ammonium mg/L as 184.1 80154 Suspended sediment concentration TSS mg/L 305.4
Group Name Type Description Unit Runoff runoff time-varying Area normalized streamflow from USGS m/y Meteorological forcings pr time-varying Daily total precipitation mm/day sph time-varying Specific humidity unitless srad time-varying Surface downwelling solar radiation tmmn time-varying Daily minimum 2-meter air temperature F tmmx time-varying Daily maximum 2-meter air temperature F pet time-varying Reference grass evapotranspiration mm/day etr time-varying Reference alfalfa evapotranspiration mm/day Rainfall chemistry pH time-varying Logarithm of the H ion activity unitless Cond time-varying Electrical conductivity of water time-varying Ca ion concentration mg/L time-varying Mg ion concentration mg/L time-varying K ion concentration mg/L time-varying Na ion concentration mg/L time-varying concentration mg/L time-varying concentration mg/L time-varying Cl ion concentration mg/L time-varying concentration mg/L distNTN time-varying The distance to the nearest NTN sampling site km Vegetation indices LAI time-varying Leaf area index of vegetation FAPAR time-varying Fraction of absorbed photosynthetically active radiation unitless NPP time-varying Net primary production Time variables datenum time-varying The number of days relative to January 1, 2000 unitless sinT time-varying Sine of datenum unitless cosT time-varying Cosine of datenum unitless Basic characteristics HYDRO_DISTURB_INDX static Hydrologic “disturbance index” score, based on 7 variables: 1) MAJ_DDENS_2009, 2) WATER_WITHDR, 3) change in dam storage 1950-2009, 4) CANALS_PCT, 5) RAW_DIS_NEAREST_MAJ_NPDES, 6) ROADS_KM_SQ_KM, and 7) FRAGUN_BASIN unitless BAS_COMPACTNESS static Watershed compactness ratio, = area/ * 100; higher number = more compact shape unitless DRAIN_SQKM static Watershed drainage area, sq km, as delineated in our basin boundary Geology GEOL_REEDBUSH_DOM static Dominant (highest percent of area) geology, derived from a simplified version of Reed & Bush (2001) - Generalized Geologic Map of the Conterminous United States unitless GEOL_REEDBUSH_DOM_PCT static Percentage of the watershed covered by the dominant geology type percentage Hydrologic characteristics STREAMS_K_S_KM static Stream density, km of streams per watershed sq km, from NHD 100k streams STRAHLER_MAX static Maximum Strahler stream order in the watershed, from NHDPlus unitless MAINSTEM_SINUOUSITY static Sinuosity of mainstem stream line, from our delineation of mainstem stream lines. Defined as curvilinear length of the mainstem stream line dividedby the straight-line distance between the end points of the line. unitless BFI_AVE static Base Flow Index (BFI). The BFI is a ratio of base flow to total streamflow, expressed as a percentage and ranging from 0 to 100. Base flow is the sustained, slowly varying component of streamflow, usually attributed to ground-water discharge to a stream. percentage CONTACT static Subsurface flow contact time index days PCT_1ST_ORDER static Percent of stream lengths in the watershed which are first-order streams (Strahler order); from NHDPlus & percentage percentage PCT_2ND_ORDER static Percent of stream lengths in the watershed which are second-order streams (Strahler order); from NHDPlus & percentage percentage PCT_3RD_ORDER static Percent of stream lengths in the watershed which are third-order streams (Strahler order); from NHDPlus & percentage percentage PCT_4TH_ORDER static Percent of stream lengths in the watershed which are fourth-order streams (Strahler order); from NHDPlus & percentage percentage PCT_5TH_ORDER static Percent of stream lengths in the watershed which are fifth-order streams (Strahler order); from NHDPlus & percentage percentage PCT_6TH_ORDER_OR_MORE static Percent of stream lengths in the watershed which are sixth or greater-order streams (Strahler order); from NHDPlus & percentage percentage
Group Name Type Description Unit Historical and current dams information DDENS_2009 static Dam density; number per 100 km sq number of dams/100 STOR_NOR_2009 static Dam storage in watershed (“NORMAL_STORAGE”); megaliters total storage per sq km (1 megalitres = 1,000,000 liters = 1,000 cubic meters) megaliters/ NPDES NPDES_MAJ_DENS static Density of NPDES (National Pollutant Discharge Elimination System) “major” point locations in the watershed; number per 100 km sq. Major locations are defined by an EPA-assigned major flag. From the download of NPDES national database summer 2006. number of sites/100 Percentages of land cover 2006 in the watershed and lanscape DEVNLCD06 static Watershed percent “developed” (urban), 2006 era (2001 for AK-HI-PR). Sum of classes 21, 22, 23, and 24. percentage FORESTNLCD06 static Watershed percent “forest”, 2006 era (2001 for AK-HI-PR). Sum of classes 41, 42, and 43. percentage PLANTNLCD06 static Watershed percent “planted/cultivated” (agriculture), 2006 era (2001 for AK-HI-PR). Sum of classes 81 and 82. percentage WATERNLCD06 static Watershed percent Open Water (class 11) percentage WOODYWETNLCD06 static Watershed percent Woody Wetlands (class 90) percentage EMERGWETNLCD06 static Watershed percent Emergent Herbaceous Wetlands (class 95) percentage Nitrogen and phosphorus application rate in the watershed NITR_APP_KG_SQKM static Estimate of nitrogen from fertilizer and manure, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed. kg/ PHOS_APP_KG_SQKM static Estimate of nitrogen from fertilizer and manure, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed. kg/ Pesticide PESTAPP_KG_SQKM static Estimate of agricultural pesticide application (219 types), kg/sq km, from Census of Ag 1997, based on county-wide sales and percent agricultural land cover in the watershed kg/ Regions ECO2_BAS_DOM static Dominant (highest % of the area) Level II ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. unitless ECO3_BAS_DOM static Dominant (highest % of the area) Level III ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. Level III ecoregion (1-84) NUTR_BAS_DOM static Dominant (highest % of the area) nutrient ecoregion within the watershed. See X_Region_Names sheet for crosswalk to name. Nutrient ecoregion (1-14) HLR_BAS_DOM_100M static Dominant (highest % of the area) Hydrologic Landscape Region within the watershed. See X_Region_Names sheet for crosswalk to name. HLR region (1-20) PNV_BAS_DOM static Dominant (highest % of the area) Potential Natural Vegetation (PNV) within the watershed. See X_Region_Names sheet for crosswalk to name. PNV type (1-63) Soil AWCAVE static Average value for the range of available water capacity for the soil layer or horizon (inches of water per inch of soil depth) unitless PERMAVE static Average permeability (inches/hour) inches/hour BDAVE static Average value of bulk density (grams per cubic centimeter) grams per cubic centimeter OMAVE static Average value of organic matter content (percent by weight) percentage WTDEPAVE static Average value of depth to seasonally high water table (feet) feet ROCKDEPAVE static Average value of total soil thickness examined (inches) inches CLAYAVE static Average value of clay content (percentage) percentage SILTAVE static Average value of silt content (percentage) percentage KFACT_UP static Average K-factor value for the uppermost soil horizon in each soil component. K-factor is an erodibility factor which quantifies the susceptibility of soil particles to detachment and movement by water. The K-factor is used in the Universal Soil Loss Equation (USLE) to estimate soil loss by water. Higher values of the K-factor indicate greater potential for erosion unitless RFACT static Rainfall and Runoff factor (“R factor” of Universal Soil Loss Equation); average annual value for the period 1971-2000. 100s ft-tonf in/h/ac/yr Topographic characteristics ELEV_MEAN_M_BASIN static Mean watershed elevation (meters) from 100m National Elevation Dataset m SLOPE_PCT static Mean watershed slope, percent. Derived from 100m resolution National Elevation Dataset, so slope values may differ from those calculated from data of other resolutions. percentage ASPECT_DEGREES static Mean watershed aspect, degrees (degrees of the compass, 0-360). Derived from 100m resolution National Elevation Data. 0 and 360 point to north, because of the national Albers projection actual aspect may vary. degrees (0-360) Latitude and Longitude LAT_GAGE static Latitude at gage, decimal degrees decimal degrees, datum NAD83 LNG_GAGE static Longitude at gage, decimal degrees decimal degrees, datum NAD83 Snow SNOW_PCT_PRECIP static Snow percent of total precipitation estimate, mean for period 1901-2000. From McCabe and Wolock (submitted, 2008), 1km grid. percentage